# TI-DeepONet: Learnable Time Integration for Stable Long-Term Extrapolation

### v2

@@ C.3 One-dimensional Kuramuto-Sivashinsky (KS) Equation
C.4 Two-dimensional Inviscid Burgers’ Equation
Data generation:
For the 2D inviscid Burgers’ equation, we consider periodic boundary conditions on a unit square domain $(x,y)\in[0,1]^{2}$. The initial conditions $u(x,y,0)=s(x,y)$ are sampled from periodic Matérn-type Gaussian random fields with length scale $l=0.1$ and standard deviation $\sigma=0.2$. The spatial domain is discretized using a $32\times 32$ grid with $N_{x}=N_{y}=32$ points. The PDE is solved using a fourth-order Runge-Kutta scheme with timestep $\Delta t=10^{-4}$ up to final time $t=1.0$. Solutions are saved at intervals of $\Delta t_{\text{save}}=0.01$, resulting in 101 temporal snapshots. The kinematic viscosity is set to $\nu=10^{-4}$, which is sufficiently small to approximate inviscid behavior while maintaining numerical stability. We generate a total of 1000 samples, each consisting of the initial condition and the corresponding spatiotemporal solution field $u(x,y,t)$ of size $101\times 64\times 64$.
Data preparation:
The dataset is partitioned using a train-test split ratio of 0.8, yielding $N_{\text{train}}=800$ training samples and $N_{\text{test}}=200$ testing samples. To evaluate extrapolation capabilities, we train only on the first third of the temporal domain, i.e., $t_{\text{train}}\in[0,0.33]$ (33 out of 101 time steps), and assess performance over the entire domain $t\in[0,1.0]$. This represents the most challenging extrapolation scenario in our study, requiring $3\times$ temporal extrapolation with no exposure to shock formation during training.
For full rollout methods, our aim is to learn a solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ that maps initial conditions to the full solution: $u_{0}(x,y)\rightarrow u(t,x,y)$. The trunk network receives $(x,y,t_{\text{train}})$ coordinates as input, while the branch network processes the initial conditions. For FNO FR, the 3D Fourier transform is performed along all three dimensions $(x,y,t_{\text{train}})$, capturing the full spatiotemporal frequency content within the training domain. In contrast, autoregressive and TI-based methods use only spatial coordinates $(x,y)$ for the trunk network input. Their branch networks receive stacked solution states $u^{0}(x,y),u^{1}(x,y),\ldots,u^{33}(x,y)]$ from the training temporal window. FNO AR employs a 2D Fourier transform along the spatial dimensions $(x,y)$ only, consistent with its autoregressive nature to learn a one-step mapping between pairs of consecutive solution states $(u^{i}(x,y),u^{i+1}(x,y))$. Note that for the autoregressive/TI-based variants, explicit temporal information is absent and instead is assimilated in the batch dimension.
Training architecture
Tables LABEL:tab:2d_burgers_don-architecture and A10 presents the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, batch size and Fourier modes (if applicable). All DeepONet variants employ Fourier feature encoding in the trunk network. For the full rollout DeepONet, the spatial and temporal coordinates $(x,y,t)$ are augmented with sinusoidal features computed as $\sin(\pi kx),\cos(\pi kx),\sin(\pi ky),\cos(\pi ky),\sin(\pi kt),\cos(\pi kt)]$ for $k\in\{1,2,\ldots,10\}$, expanding the input dimension from 3 to 63. For the autoregressive and TI-based variants, only the spatial coordinates $(x,y)$ are augmented with Fourier features $\sin(\pi kx),\cos(\pi kx),\sin(\pi ky),\cos(\pi ky)]$, expanding the input dimension from 2 to 42. In this case, for the TI(L)-DeepONet architecture, we employ a similar combination of 2D convolutional blocks and MLP layers as used in the branch network to learn the RK4 slope coefficients. The convolutional blocks utilize the same combination of Conv2D and pooling layers as the branch, with the only difference being that the Conv2D layers have a reduced number of filters (32 instead of 64). This is followed by a flattening layer, after which the intermediate representation is passed through two fully connected hidden layers and finally to an output layer with 4 neurons that predicts the RK4 slope coefficients.
The Conv2D layers use a ReLU activation function, the fully connected hidden layers use tanh, and the output layer applies a softmax activation to ensure that the predicted slope coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 steps. The auxiliary network responsible for predicting the learnable RK4 slope coefficients was trained with a similar exponential decay schedule, but with a higher initial learning rate of $2\times 10^{-3}$. For both FNO AR and FNO FR variants, we employ a similar exponential decay schedule as the DeepONet variants albeit with a decay factor of 0.96 every 2000 steps.
Table A9: 2D Inviscid Burgers’ equation: Neural network architecture and training settings across different methods.
@@ C.4 Two-dimensional Inviscid Burgers’ Equation
Figure A2: 2D Inviscid Burgers’ equation: Solution field for a representative sample, shown across both the training regime ($t\in[0,0.33]$) and the extrapolation regime ($t\in[0.33,1]$), for all frameworks. Corresponding error plots are presented in Figure 7.

### v3

@@ B.3 One-dimensional Kuramuto-Sivashinsky (KS) Equation
B.4 Two-dimensional Burgers’ Equation
Table A8: 2D Burgers’ Equation: Neural network architecture and training settings across different methods. Note that each epoch corresponds to a single gradient descent update over a randomly sampled batch, equivalent to one stochastic gradient descent iteration.
@@ B.4 Two-dimensional Burgers’ Equation
Figure A2: 2D Burgers’ Equation: Solution field for a representative sample, shown across both the training regime ($t\in[0,0.33]$) and the extrapolation regime ($t\in[0.33,1]$), for all frameworks. Corresponding error plots are presented in Fig. 12.
B.5 Two-dimensional Scalar Rotation Advection-Diffusion
Table A10: 2D Scalar Rotation Advection-Diffusion: Neural network architecture and training settings across different methods. Note that each epoch corresponds to a single gradient descent update over a randomly sampled batch, equivalent to one stochastic gradient descent iteration.
Method Branch Net Trunk Net Activation Epochs Batch Size
[…]
Table A11: 2D Scalar Rotation Advection-Diffusion: Neural network architecture and training settings across different FNO-based methods.
Method Fourier Blocks Modes Hidden Dim Activation Epochs Batch Size
FNO AR 6 12 64 GELU 1000 64
FNO FR 6 12 64 GELU 5000 32
Figure A3: 2D Rotation Advection-Diffusion equation: Solution field for a representative sample, shown across both the training regime ($t\in[0,0.27]$) and the extrapolation regime ($t\in[0.27,0.67]$), for all frameworks. Corresponding error plots are presented in Fig. 16.
B.6 Three-dimensional Heat Conduction
Table A12: 3D Heat Conduction: Neural network architecture and training settings across different methods. Note that each epoch corresponds to a single gradient descent update over a randomly sampled batch, equivalent to one stochastic gradient descent iteration.
Method Branch Net Trunk Net Activation Epochs Batch Size
[…]
Figure A4: 3D Heat Conduction: Solution field for a representative sample, shown across both the training regime ($t\in[0,0.33]$) and the extrapolation regime ($t\in[0.33,1]$), for all frameworks. Corresponding error plots are presented in Fig. 19. The contours are presented along the XY slicing plane (see Fig. 18 for details regarding the location of the XY slice).
Figure A5: 3D Heat Conduction: Solution field for a representative sample, shown across both the training regime ($t\in[0,0.33]$) and the extrapolation regime ($t\in[0.33,1]$), for all frameworks. Corresponding error plots are presented in Fig. 20. The contours are presented along the YZ slicing plane (see Fig. 18 for details regarding the location of the YZ slice).
Figure A6: 3D Heat Conduction: Solution field for a representative sample, shown across both the training regime ($t\in[0,0.33]$) and the extrapolation regime ($t\in[0.33,1]$), for all frameworks. Corresponding error plots are presented in Fig. 21. The contours are presented along the XZ slicing plane (see Fig. 18 for details regarding the location of the XZ slice).